If GMAT scores for applicants at Oxnard Graduate School of Business are N(500, 50), then the top 5 percent of the applicants would have a score of at least (choose the nearest integer)

Multiple Choice

  • 601

  • 582

  • 608

  • 575
Answer #1

Solution:-

Given that,

mean = = 500

standard deviation = =50

Using standard normal table,

P(Z > z) = 5%

= 1 - P(Z < z) = 0.05

= P(Z < z) = 1 - 0.05

= P(Z < z ) = 0.95

= P(Z <1.65 ) = 0.95

z =1.65 ( using z table )

Using z-score formula,

x = z * +

x = 1.65 * 50+500

x = 582